an arrow of mass 0.018 kg is shot upwards at an angle from the ground, reaching a maximum height of 25 m with a speed of 62 m/s. Neglecting air resistance, with what speed will the arrow hit the ground?? please show all work rly need help
1
Expert's answer
2019-12-19T11:36:51-0500
"v_x=62\frac{m}{s}"

"v_y=\sqrt{2gh}=\sqrt{2(9.8)(25)}=22.1\frac{m}{s}"

Final speed:


"v=\sqrt{(v_x)^2+(v_y)^2}=\sqrt{(62)^2+(22.1)^2}=65.8\frac{m}{s}"